OpenStudy (anonymous):

6 + 7 * log2 x = 14 with 2 being the base of the log

OpenStudy (anonymous):

subtract 6 from both sides\[6 - 6 + 7*\log2 x = 14 - 6\] Divide 7 on both sides\[(7 * \log2 x )/7= 8/7\] This gives \[\log 2 x = 8/7\] If 2 is the base of the log, than x will equal the base of the log to an exponent.\[2^(8/7) = x\]